Consider the linear equation 2a + 3b = 4


Is (a; b) = (1; 1


2


) a solution to the equation? Motivate your answer.

1
Expert's answer
2022-03-07T08:12:03-0500

If "(a, b)=(1,\dfrac{1}{2})," then substitute


"2(1) + 3(\dfrac{1}{2}) = 4"

"\dfrac{5}{2}=4, False"

Therefore "(a, b)=(1, \dfrac{1}{2})" is not a solution to the equation "2a+3b=1."


If "(a, b)=(\dfrac{1}{2}, 1)," then substitute


"2(\dfrac{1}{2}) + 3(1) = 4"

"4=4, True"

Therefore "(a, b)=(\dfrac{1}{2}, 1)" is a solution to the equation "2a+3b=1."